OpenStudy (anisah.21a):

13=3log(x)+25 log(48x^3)-log(8x^2)=4 4ln(x-11)-2=30 solve for x on all three

OpenStudy (deadly_roses):

1. = 1/e4 2. = e4/6 3. = 11 + e8 this is what i got...im not sure if this is the format your looking for...